# Probing dark-matter effects with gravitational waves using the parameterized post-Einsteinian framework

### v1

@@ I Introduction
Since presence of dark matter in an IMRI can produce a measurable GW dephasing from analogous vacuum systems, it is useful to know how precisely this dephasing needs to be modeled to be detectable and have minimal bias in the inferred GW parameters. While using the true signal is the optimal choice for matched-filtering-based data analysis, there are more model-agnostic frameworks, such as the parametrized post-Einsteinian (ppE) approach [37, 38, 39], which aim to capture deviations from vacuum waveforms in general relativity (GR). The ppE framework has been shown to detect deviations from GR in simulation studies [40, 41, 42] and also has been applied to the existing GW events to constrain such deviations [10, 11, 17, 18, 19]. Parameterized waveforms similar to those used in the ppE framework have also been applied to study effects of the astrophysical environment in IMRIs, in particular those induced by dark matter (see in [43, 44]). However, neither of the signal models used in [43, 44] were as complex as the phase in the dynamic case in [36], so it has not been established how well the ppE framework can capture effects induced by frequency-domain signals with a more complicated morphology.
@@ V Results of the ppE tests
Given these very large biases on the chirp-mass parameter estimation, it is unlikely that a vacuum template would be capable of detecting an IMRI within a DM spike using a matched-filtering search. Thus, such a study implicitly assumes that the signal was detected by some other means (such as a coherent excess-power search) and was analyzed with vacuum templates.
@@
Figure 6 shows the errors on the chirp mass and ppE $\beta$ with the updated fiducial values. For the chirp mass, the statistical error remains almost unaffected from the results in Q2, while the systematic error has been reduced by a few orders of magnitude. For ppE parameter, the fractional systematic error is $\mathcal{O}(10^{-11})$ in most cases; however, the fractional statistical error is $\mathcal{O}(10^{-3})$. The fractional statistical error is roughly consistent with the ratio of the absolution statistical and systematic errors in Q2. It is not precisely the same, however, because the partial derivative $\partial\tilde{h}/\partial\mathcal{M}_{z}$ in the inverse Fisher matrix now contains contributions from terms proportional to $\beta$, which is assumed to be nonzero. This will introduce additional correlations between $\mathcal{M}_{z}$ and $\beta$, which would vanish when the fiducial value of $\beta$ is set to be zero, as in Q2. Nevertheless, the systematic errors are now much smaller than the statistical errors for both parameters. This shows that, in principle, the ppE template is able to reduce the bias in vacuum parameters and to capture the effects of a dynamical DM distribution with a nonzero fiducial value of $\beta$.
@@ VI Conclusions
The systematic errors on the vacuum parameters depend sensitively on the choice of the effective PN order; specifically, if the value is slightly off from the optimized one, the error goes up by several orders of magnitude. To run the ppE framework with a fixed exponent $b$, it would be necessary to perform the analysis for a large number of finely sampled $b$ values. This may not be feasible, so it is more natural to treat the ppE parameter $b$ as an undetermined model parameter that has its own systematic and statistical errors. Using the same optimized value from the previous test caused the systematic errors to exceed the statistical errors, and attempts at further optimization of the fiducial parameters did not significantly improve the results. We conclude that although the ppE template can, in principle, remove the bias in the parameter estimation for GWs from binary black holes with a DM halo, it may be difficult in practice to do so using the the simplest ppE framework.

### v2

@@ I Introduction
Since presence of dark matter in an IMRI can produce a measurable GW dephasing from analogous vacuum systems, it is useful to know how precisely this dephasing needs to be modeled to be detectable and have minimal bias in the inferred GW parameters. While using the true signal is the optimal choice for matched-filtering-based data analysis, there are more model-agnostic frameworks, such as the parametrized post-Einsteinian (ppE) approach [37, 38, 39], which aim to capture deviations from vacuum waveforms in general relativity (GR). The ppE framework has been shown to detect deviations from GR in simulation studies [40, 41, 42] and also has been applied to the existing GW events to constrain such deviations [10, 11, 17, 18, 19]. Parameterized waveforms similar to those used in the ppE framework have also been applied to study effects of the astrophysical environment in IMRIs, in particular those induced by dark matter (see in [43, 44, 45]). However, neither of the signal models used in [43, 45] were as complex as the phase in the dynamic case in [36], so it has not been established how well the ppE framework can capture effects induced by frequency-domain signals with a more complicated morphology.
@@ V Results of the ppE tests
The results above show that if a GW signal from an IMRI within a DM spike is detected through a matched-filtering search against vacuum templates, there will be large biases on the chirp mass. Given the results in Fig. 6 of [36], however, it will likely be challenging to detect such systems with vacuum templates. Thus, if a vacuum template is not capable of detecting the signal, then we would assume that the signal was detected by some other means (such as a coherent excess-power search) and was analyzed with vacuum templates.
@@
Figure 6 shows the errors on the chirp mass and ppE $\beta$ with the updated fiducial values. For the chirp mass, the statistical error remains almost unaffected from the results in Q2, while the systematic error has been reduced by a few orders of magnitude. For ppE parameter, the fractional systematic error is $\mathcal{O}(10^{-10})$ in most cases; however, the fractional statistical error is $\mathcal{O}(10^{-4})$. The fractional statistical error is roughly consistent with the ratio of the absolute statistical and systematic errors in Q2. It is not precisely the same, however, because the partial derivative $\partial\tilde{h}/\partial\mathcal{M}_{z}$ in the inverse Fisher matrix now contains contributions from terms proportional to $\beta$, which is assumed to be nonzero. This will introduce additional correlations between $\mathcal{M}_{z}$ and $\beta$, which would vanish when the fiducial value of $\beta$ is set to be zero, as in Q2. Nevertheless, the systematic errors are now much smaller than the statistical errors for both parameters. This shows that, in principle, the ppE template is able to reduce the bias in vacuum parameters and to capture the effects of a dynamical DM distribution with a nonzero fiducial value of $\beta$.
@@ VI Conclusions
The systematic errors on the vacuum parameters depend sensitively on the choice of the effective PN order; specifically, if the value is slightly off from the optimized one, the error goes up by several orders of magnitude. To run the ppE framework with a fixed exponent $b$, it would be necessary to perform the analysis for a large number of finely sampled $b$ values. This may not be feasible, so it is more natural to treat the ppE parameter $b$ as an undetermined model parameter that has its own systematic and statistical errors. Using the same optimized value from the previous test caused the systematic errors to exceed the statistical errors, and attempts at further optimization of the fiducial parameters did not significantly improve the results. We conclude that although the ppE template can, in principle, remove the bias in the parameter estimation for GWs from binary black holes with a DM halo, it may be difficult in practice to do so using the simplest ppE framework.
@@
Finally, the intent of the original ppE framework is to test for theory-agnostic deviations from GR that arise at fixed PN orders, most frequently at half-integer values. Our results have shown that a nonzero ppE parameter $\beta$ significantly reduces the systematic error in our Fisher-based parameter-estimation studies, which suggests that waveforms with nonzero $\beta$ should be used to analyze the GW signals from such systems. In the context of the ppE framework, a nonzero $\beta$ is associated with a deviation from GR, which might suggest that an environmental effect from a dynamic DM spike instead could be interpreted as a deviation from GR. However, the systematic error depended sensitively on the ppE exponent parameter $b$, which took on a continuum of values rather than those associated with a particular half-integer PN order. The fact the best value of $b$ is not one of these discrete values is a likely indication that the origin of the nonzero $\beta$ is an environmental effect rather than a beyond-GR effect. However, it may be the case that in practical ppE studies, the tests will be performed only a handful of half-integer PN orders. As a future study, it would also be interesting to investigate how well parameter estimation with ppE waveforms with these discrete ppE exponents $b$ perform when the waveform of an IMRI with a dynamic DM spike is used as the injected signal. Such studies would determine whether or not the environmental effects arising from a dynamic DM spike could mimic those effects more commonly associated with gravitational theories beyond GR.